# Geometric Autoencoder Priors for Bayesian Inversion: Learn First Observe Later

### v1

@@ Abstract
Uncertainty Quantification (UQ) is paramount for inference in engineering applications. A common inference task is to recover full-field information of physical systems from a small number of noisy observations, a usually highly ill-posed problem. Critically, engineering systems often have complicated and variable geometries prohibiting the use of standard Bayesian UQ. In this work, we introduce Geometric Autoencoders for Bayesian Inversion (GABI), a framework for learning geometry-aware generative models of physical responses that serve as highly informative geometry-conditioned priors for Bayesian inversion. Following a “learn first, observe later” paradigm, GABI distills information from large datasets of systems with varying geometries, without requiring knowledge of governing PDEs, boundary conditions, or observation processes, into a rich latent prior. At inference time, this prior is seamlessly combined with the likelihood of the specific observation process, yielding a geometry-adapted posterior distribution. Our proposed framework is architecture agnostic. A creative use of Approximate Bayesian Computation (ABC) sampling yields an efficient implementation that utilizes modern GPU hardware. We test our method on: steady-state heat over rectangular domains; Reynold-Averaged Navier-Stokes (RANS) flow around airfoils; Helmholtz resonance and source localization on 3D car bodies; RANS airflow over terrain. We find: the predictive accuracy to be comparable to deterministic supervised learning approaches in the restricted setting where supervised learning is applicable; UQ to be well calibrated and robust on challenging problems with complex geometries. The method provides a flexible geometry-aware train-once-use-anywhere foundation model which is independent of any particular observation process.
@@ 1 Introduction
Many important problems in engineering deal with spatially varying quantities that must be inferred. The task of inferring full-field information from sparse noisy measurement is a type of inverse problem (Kaipio & Somersalo, 2005); these often need to be regularized due to their inherent ill-posedness (Calvetti & Reichel, 2003). Statistically interpretable regularization can be achieved through the Bayesian paradigm (Stuart, 2010), where we combine a likelihood function characterizing data fit for a particular observation process, and a prior which characterizes assumptions about the data-generating process. Bayesian inference allows for robust and principled uncertainty quantification, where we obtain a distribution on the estimated quantity of interest given data from a particular physical system. The main challenge in enacting a Bayesian program is prior specification; often, not much can be said about the prior which often leads to diffuse and vague posteriors with relatively poor predictive accuracy when compared to deterministic methods. To alleviate this vagueness, one possible strategy is to learn a suitable prior from a dataset of related but distinct physical setups where more data may be available. This is explored in general-purpose AI applications (Shwartz-Ziv et al., 2022; Feng et al., 2023; Boys et al., 2024) and physical applications (Meng et al., 2022; Akyildiz et al., 2025; Patel & Oberai, 2021; Patel et al., 2022; Laloy et al., 2018; 2017). However, in engineering, geometry plays a crucial role in determining the dynamics arising for a physical system. Hence, for truly general purpose, practical methodology, learned priors must incorporate knowledge about the geometry of the problem at hand. But changing geometries hinder the direct learning of a prior over fields as the probability spaces involved are tied to the geometry of each individual physical system, prohibiting a direct Bayesian treatment.
In this work we propose a methodology, using graph autoencoders, to learn highly informative priors as geometry-aware generative models for use in Bayesian inversion tasks. We also show how to effectively solve the Bayesian inverse problems resulting from these priors through a pushfoward model. Importantly, the proposed scheme decouples the task of “learning” and the task of ”inferring from observations”, resulting in train-once use everywhere approach which is independent of the particular observation process for solving inverse problems, a trait not shared by related supervised learning approaches. We make the following contributions:
@@
    We propose a methodology to encode geometries and physical fields into a learned latent prior; without specifying the governing PDE or boundary conditions.
@@
    The proposed methodology is naturally extended to the Bayesian estimation of other important information about the observation process such as learning the observational noise.
@@ 1.1 Setup
Let $u_{n}\in\mathcal{U}_{n}:=\mathcal{U}(\mathcal{M}_{n};\mathbb{R}^{d_{u}})$ be a (possibly vector valued) function representing the full field state of a physical system on the bounded domain $\mathcal{M}_{n}\subset\mathbb{R}^{d}$. This domain may be a manifold as we will see later. In this work, $u_{n}$ will be: heat on a surface, fluid flow in a domain, and resonance patterns on an object. Here the index $n$ refers to the different physical systems which have possibly different geometries, forcing, and boundary conditions. In numerically representing these continuous systems we obtain discretized node values $\mathbf{u}_{n}\in\mathsf{U}_{n}:=\mathsf{U}(\mathsf{M}_{n};\mathbb{R}^{d_{u}})$ which live on $\mathsf{M}_{n}=(\mathsf{V}_{n},\mathcal{E}_{n})$ where $\mathsf{V}_{n}=\{x_{1},\ldots,x_{d_{n}}\}\subset\mathcal{M}_{n}$ is a set of vertices and $\mathcal{E}_{n}$ is a set of edges connecting the vertices. As such, we represent $\mathsf{M}_{n}$ as an undirected graph describing the physical geometries on which the physics is taking place, i.e. a computational mesh. We assume to have a dataset $\mathcal{D}=\{\mathbf{u}_{n},\mathsf{M}_{n}\}_{n=1}^{N}$  [^1] A dataset with multiple full-fields per geometry $\mathsf{M}_{n}$, as $\{\{\mathbf{u}_{n,i}\}_{i=1}^{N_{n}},\mathsf{M}_{n}\}_{n=1}^{N}$, can viewed as a dataset $\{\mathbf{u}_{k},\mathsf{M}_{k}\}_{k=1}^{K}$ indexed by a dummy variable $k$ such that $k$ maps to $(i,n)$.. This dataset can be constructed from simulation data, as is the case in this paper, or from full-field data from experiment such as through particle image velocimetry, digital image correlation, magnetic resonance, electrical impedance tomography, infrared thermography etc.
@@
First, in the proposed methodology we learn a $\mathsf{M}_{n}$-dependent prior generative model for $\mathbf{u}_{n}$. Second, using this learned highly informative prior and sparse noisy observations, $\mathbf{y}_{o}\in\mathbb{R}^{d_{o}}$  [^2] When discussing observations which are associated to a geometry not in the dataset $\mathcal{D}$ we change the index from $n$ to $o$ for clarity., of a new physical system on a new geometry, $\mathsf{M}_{o}$, we solve the Bayesian inverse problem to obtain the posterior distribution $p_{\mathbf{u}_{o}|\mathbf{y}_{o}}$. We assume to know the discretized geometry $\mathsf{M}_{o}$ associated to $\mathbf{y}_{o}$. In this paper, the observations and discretized solution field are related by:
$\mathbf{y}_{o}=\mathbf{H}_{o}\mathbf{u}_{o}+\bm{\xi}_{o},\quad\bm{\xi}_{o}\sim\mathcal{N}(\mathbf{0},\sigma^{2}{\mathbf{I}}_{o});$ (1)
$\mathbf{H}_{o}:\mathbb{R}^{d_{o}}\times\mathbb{R}^{d_{u}}\rightarrow\mathbb{R}^{o}$ is the observation operator selecting the nodes to be observed. Other observation models can be considered with ease.
@@ 1.2.1 Direct Map Inversion
Many interesting approaches have been recently proposed which directly learn an amortized map from observable to solution/parameter. These can be subdivided into deterministic methods Bayesian/probabilistic approaches. Most related to the setting considered in this paper, is the work in Duthé et al. (2025). Here the authors consider a graph-ML approach to learning a deterministic map from observable $\mathbf{y}_{o}$ to the underlying vector field $\mathbf{u}_{o}$. The supervised methodology used in the comparisons in this work take inspiration from this work. Another strongly related methodology can be found in Arridge et al. (2019) although here the problem of changing geometry is not directly tackled. Other noteworthy works in this area include (Adler & Öktem, 2017; Dittmer et al., 2020) Interesting works look at probabilistic schemes for constructing probabilistic maps from observable to quantity of interest (Kaltenbach et al., 2023; Baptista et al., 2024). These works do not focus on learning priors. Other works in this area include (Adler et al., 2022; Ardizzone et al., 2018).
@@ 1.2.2 Forward and Inverse Problems on Graphs
Graphs are handled in machine learning frameworks through geometric deep learning (Bronstein et al., 2021; Wu et al., 2022; Kipf, 2016). Several important works consider inverse problems where the quantity inferred lives of a graph description of the geometry. In Garcia Trillos & Sanz-Alonso (2018) the authors prove the convergence of the graph posterior to the continuum posterior subject to some conditions. In Povala et al. (2022) the adjacency structure of a mesh is leveraged to accelerate variational inference inversion. Important works consider solving the PDE forward problem through physics informed graph machine learning(Seo & Liu, 2019; Valencia et al., 2025; Wandel et al., 2020) Used in our comparisons are linear inverse problems with Gaussian process priors, this leads to graph Gaussian process methods (Borovitskiy et al., 2021; Mostowsky et al., 2024). We find it important to mention important works in the physics informed literature, where assuming complete or partial knowledge PDEs, boundary conditions, forcing etc, one can use physics residuals to learn forward and inverse maps for PDEs (Raissi et al., 2019; Gao et al., 2022; Vadeboncoeur et al., 2023; Xu et al., 2023; Rixner & Koutsourelakis, 2021). However, in this work we focus solely on methods which do not assume knowledge of the governing PDE, boundary conditions, or forcing and are entirely data-driven in their approach.
@@ 1.2.3 Statistical Autoencoders
Related to inverse problem are latent variable representations of graph data structures. In this category the seminal work of the graph VAE (Kipf & Welling, 2016; Nazari et al., 2023). We refer the reader to Appendix C for a more elaborate discussion on the relationship between VAEs and our proposed framework. Related to this are the works (Mylonas et al., 2021; Liu et al., 2019) The proposed work centres around a class of autoencoders dubbed statistical autoencoders, of which there are many variants (Zhao et al., 2019; Kolouri et al., 2019; Turinici, 2021; Tolstikhin et al., 2018). We focus in this work on the MMD (Gretton et al., 2012) variant of the autoencoder. Many other variants could be considered. Another relevant class of autoencoders are conditional autoencoders (Sohn et al., 2015), we will make use of related ideas by conditioning encoder and decoder on geometry.
@@ 1.3 Notation
We denote by $\mathbb{P}_{z}$ a probability measure on the measurable space $(\mathcal{Z},\mathcal{S})$, and by $p_{z}$ the associated density with respect to Lebesgue measure when it exists. When $\mathbb{P},\mathbb{Q}$ are empirical measures, i.e. a mixture of Dirac measures, we will sometimes refer to these measures as the “densities” $p,q$ to lighten notation. Superscripts in parentheses denote realizations of random variables. For a function $g:\mathcal{Z}\rightarrow\mathcal{U}$, the pushforward $g_{\#}p_{z}$ means $g_{\#}\mathbb{P}_{z}(A)=\mathbb{P}_{z}(g^{-1}(A))$ for all measurable sets $A$. We note a sample $u^{(n)}\sim g_{\#}p_{z}$ is obtained as $u^{(i)}=g(z^{(i)}),z^{(i)}\sim p_{z}$.
@@ 2.1 Posterior from a Pushforward Prior
All proofs can be found in Appendix A. This result will be explicitly related to the geometric inference setting in Theorem 2.2, showing that with a geometry agnostic latent space, constructed via a geometry aware encoder/decoder, we can solve inverse problems with a single common latent representation relevant to any number of geometries and learn highly informative priors across geometries.
@@ 2.2 Geometric Autoencoder Priors for Bayesian Inversion (GABI)
We describe the 2 step process of learning a geometry-aware prior and how to use this for solving inverse problems.
@@
These two maps imply a geometry conditioned autoencoder. We use $\widehat{\mathbb{E}}_{\mathcal{D}}[\,\cdot\,]$ as a shorthand for $\frac{1}{N}\sum_{(\mathbf{u}_{n},\mathsf{M}_{n})\in\mathcal{D}}[\,\cdot\,]$. Set the parameters of the auto-encoder to
@@
where $p^{\theta}_{\mathbf{z}}=\widehat{\mathbb{E}}_{\mathcal{D}}\,[\delta_{E^{\theta}_{n}(\mathbf{u}_{n})}]$ is the empirical distribution of encoded solutions, $q_{\mathbf{z}}=\mathcal{N}(\bf 0,\mathbf{I})$, and $\mathsf{d}:\mathcal{P}(\mathsf{Z})\times\mathcal{P}(\mathsf{Z})\rightarrow[0,\infty)$ may be any suitable divergence that is well-posed for comparing empirical measures: maximum mean discrepancy (MMD), energy distance (ED), Wasserstein, or Sliced-Wasserstein distances etc. In this work, we focus on MMD.
@@
$\mathbf{y}_{o}=\mathbf{H}_{o}D^{\psi}_{o}(\mathbf{z})+\bm{\xi}_{o};\quad\bm{\xi}_{o}\sim\mathcal{N}(\mathbf{0},\sigma^{2}{\mathbf{I}}_{o}).$ (4)
We specify our prior distribution over $\mathbf{u}_{o}$ using the trained decoder and specify the likelihood associated to this decoded observation model
$\centering p_{\mathbf{u}_{o}}:=D^{\psi}_{o\,\#}q_{\mathbf{z}};\quad\quad p^{\psi}_{\mathbf{y}_{o}|\mathbf{z}}=\mathcal{N}(\mathbf{H}_{o}D^{\psi}_{o}(\mathbf{z}),\sigma^{2}\mathbf{I}_{o}).\@add@centering$ (5)
@@
Assuming the relevant measurability of $D^{\psi}_{o}=D^{\psi}(\,\cdot\,;\mathsf{M}_{o})$, the posterior
$p_{\mathbf{u}_{o}|\mathbf{y}_{o}}=D^{\psi}_{o\,\#}p_{\mathbf{z}|\mathbf{y}_{o}}.$ (7)

### v2

@@ Abstract
Uncertainty Quantification (UQ) is paramount for inference in engineering. A common inference task is to recover full-field information of physical systems from a small number of noisy observations, a usually highly ill-posed problem. Sharing information from multiple distinct yet related physical systems can alleviate this ill-possendess. Critically, engineering systems often have complicated variable geometries prohibiting the use of standard multi-system Bayesian UQ. In this work, we introduce Geometric Autoencoders for Bayesian Inversion (GABI), a framework for learning geometry-aware generative models of physical responses that serve as highly informative geometry-conditioned priors for Bayesian inversion. Following a “learn first, observe later” paradigm, GABI distills information from large datasets of systems with varying geometries, without requiring knowledge of governing PDEs, boundary conditions, or observation processes, into a rich latent prior. At inference time, this prior is seamlessly combined with the likelihood of a specific observation process, yielding a geometry-adapted posterior distribution. Our proposed framework is architecture agnostic. A creative use of Approximate Bayesian Computation (ABC) sampling yields an efficient implementation that utilizes modern GPU hardware. We test our method on: steady-state heat over rectangular domains; Reynold-Averaged Navier-Stokes (RANS) flow around airfoils; Helmholtz resonance and source localization on 3D car bodies; RANS airflow over terrain. We find: the predictive accuracy to be comparable to deterministic supervised learning approaches in the restricted setting where supervised learning is applicable; UQ to be well calibrated and robust on challenging problems with complex geometries.
@@ 1 Introduction
Many important problems in engineering deal with spatially varying quantities that must be inferred. The task of inferring full-field information from sparse noisy measurement is a type of inverse problem (Kaipio & Somersalo, 2005); these often need to be regularized due to their inherent ill-posedness (Calvetti & Reichel, 2003). Statistically interpretable regularization can be achieved through the Bayesian paradigm (Stuart, 2010), where we combine a likelihood function characterizing data fit for a particular observation process, and a prior which characterizes assumptions about the data-generating process. Bayesian inference allows for robust and principled uncertainty quantification, where we obtain a distribution on the estimated quantity of interest given data from a particular physical system. The main challenge in enacting a Bayesian program is prior specification; often, not much can be said about the prior which often leads to diffuse and vague posteriors with relatively poor predictive accuracy when compared to deterministic methods. To alleviate this vagueness, one possible strategy is to learn a suitable prior from a dataset of related but distinct physical setups, where more data may be available. This is explored in general-purpose AI applications (Shwartz-Ziv et al., 2022; Feng et al., 2023; Boys et al., 2024) and physical applications (Meng et al., 2022; Akyildiz et al., 2025; Patel & Oberai, 2021; Patel et al., 2022; Laloy et al., 2018; 2017). However, in engineering, geometry plays a crucial role in determining the dynamics arising for a physical system. Hence, for truly general purpose, practical methodology, learned priors must incorporate knowledge about the geometry of the problem at hand. But changing geometries hinder the direct learning of a prior over fields as the probability spaces involved are tied to the geometry of each individual physical system, prohibiting a direct treatment.
In this work, we propose a methodology using graph-based autoencoders to learn highly informative priors as geometry-aware generative models for use in Bayesian inversion tasks. We also show how to effectively solve the Bayesian inverse problems resulting from these priors through a pushfoward model. Importantly, the proposed scheme decouples the task of “learning” and the task of ”inferring from observations”, resulting in train-once use everywhere approach which is independent of the particular observation process for solving inverse problems, a trait not shared by related supervised learning approaches. We make the following contributions:
@@
    We propose a methodology to encode geometries and physical fields into a learned latent prior without specifying the governing PDE or boundary conditions.
@@
    The proposed methodology is naturally extended to the Bayesian estimation of other important quantities in the observation process, such as learning the observational noise.
@@ 1.1 Setup
Let $u_{n}\in\mathcal{U}_{n}:=\mathcal{U}(\mathcal{M}_{n};\mathbb{R}^{d_{u}})$ be a (possibly vector valued) function representing the full field behavior of a physical system on the bounded domain $\mathcal{M}_{n}\subset\mathbb{R}^{d}$. The space $\mathcal{U}$ is the relevant function space for the full field solution, e.g. a Sobolev space, and the geometric domain $\mathcal{M}_{n}$ may be a manifold with boundary or a boundaryless closed manifold. In this work, $u_{n}$ will be: heat on a surface, fluid flow in a domain, and resonance patterns on an object. Here, the index $n$ refers to the different physical systems which possibly have different geometries, forcing, and boundary conditions. In numerically representing these continuous systems, we obtain discretized node values $\mathbf{u}_{n}\in\mathsf{U}_{n}:=\mathsf{U}(\mathsf{M}_{n};\mathbb{R}^{d_{u}})$ which live on $\mathsf{M}_{n}=(\mathsf{V}_{n},\mathcal{E}_{n})$ where $\mathsf{V}_{n}=\{x_{1},\ldots,x_{d_{n}}\}\subset\mathcal{M}_{n}$ is a set of vertices and $\mathcal{E}_{n}$ is a set of edges connecting the vertices. As such, we represent $\mathsf{M}_{n}$ as an undirected graph describing the physical geometries on which the physics is taking place, i.e., a computational mesh. We assume to have a dataset $\mathcal{D}=\{\mathbf{u}_{n},\mathsf{M}_{n}\}_{n=1}^{N}$  [^1] A dataset with multiple full-fields per geometry $\mathsf{M}_{n}$, as $\{\{\mathbf{u}_{n,i}\}_{i=1}^{N_{n}},\mathsf{M}_{n}\}_{n=1}^{N}$, can be viewed as a dataset $\{\mathbf{u}_{k},\mathsf{M}_{k}\}_{k=1}^{K}$ indexed by a dummy variable $k$ such that $k$ maps to $(i,n)$.. This dataset can be constructed from simulation data, as is the case in this paper, or from full-field data from experiment such as through particle image velocimetry, digital image correlation, magnetic resonance, electrical impedance tomography, infrared thermography etc.
@@
First, in the proposed methodology we learn a $\mathsf{M}_{n}$-dependent prior generative model for $\mathbf{u}_{n}$. Second, using this learned highly informative prior and sparse noisy observations, $\mathbf{y}_{o}\in\mathbb{R}^{d_{o}}$  [^2] When discussing observations which are associated to a geometry not in the dataset $\mathcal{D}$ we change the index from $n$ to $o$ for clarity of exposition., of a new physical system on a new geometry, $\mathsf{M}_{o}$, we solve the Bayesian inverse problem to obtain the posterior distribution $p_{\mathbf{u}_{o}|\mathbf{y}_{o}}$. We assume to know the discretized geometry $\mathsf{M}_{o}$ associated to $\mathbf{y}_{o}$. In this paper, the observations and discretized solution field are related by:
$\mathbf{y}_{o}=\mathbf{H}_{o}\mathbf{u}_{o}+\bm{\xi}_{o},\quad\bm{\xi}_{o}\sim\mathcal{N}(\mathbf{0},\sigma^{2}{\mathbf{I}}_{o}).$ (1)
We take $\mathbf{H}_{o}:\mathbb{R}^{d_{u}}\rightarrow\mathbb{R}^{d_{o}}$ to be a matrix of zeros and ones selecting the nodes to be observed. This is assumed to be a known observation operator corresponding to sensor placements. We note that many observation models can be considered with ease, such as surface integrals for computing drag coefficients, radon transforms for tomography, as well as any number of noise models – additive or not. As our learned prior is entirely independent of the observation process, changes to the observation process do not necessitate retraining and can be incorporated “on-the-fly”.
The proposed methodology works as follows:
  • In the training stage, using a dataset of geometries and full-field solutions on these geometries, $\mathcal{D}=\{\mathbf{u}_{n},\mathsf{M}_{n}\}_{n=1}^{N}$, we learn a graph-based autoencoder which embeds the joint distribution of solutions and geometries into a latent Gaussian distribution. This will act as our highly informative geometry-aware data-driven prior.
  • In the application stage, we use this learned prior to solve inverse problems for data coming from new physical systems not in the dataset $\mathcal{D}$. That is, given sparse and noisy observations $\mathbf{y}_{o}$ of a solution $\mathbf{u}_{o}$ of a physical systems with geometry $\mathsf{M}_{o}$, we use our previously learned prior to solve the Bayesian inverse problem for the unknown full-field $\mathbf{u}_{o}$. We recover a distribution over the solution $\mathbf{u}_{o}$ by only observing $\mathbf{y}_{o}$.
@@ 1.2.1 Direct Map Inversion
Many interesting approaches have been recently proposed which directly learn an amortized map from observable to solution/parameter. These can be subdivided into deterministic methods and probabilistic - often Bayesian - approaches. Most related to the setting considered in this paper is the work in Duthé et al. (2025). Here, the authors consider a graph-ML approach to learning a deterministic map from observable $\mathbf{y}_{o}$ to the underlying vector field $\mathbf{u}_{o}$. The supervised methodology used in the comparisons in this work draws inspiration from that prior work. Another strongly related methodology can be found in Arridge et al. (2019) although here the problem of changing geometry is not directly tackled. Other noteworthy works in this area include (Adler & Öktem, 2017; Dittmer et al., 2020). Interesting works look at probabilistic schemes for constructing conditional maps from observables to quantities of interest (Kaltenbach et al., 2023; Baptista et al., 2024; Hosseini et al., 2025). Other works in this area include (Adler et al., 2022; Ardizzone et al., 2018).
@@ 1.2.2 Forward and Inverse Problems on Graphs
Graphs are handled in machine learning frameworks through geometric deep learning (Bronstein et al., 2021; Wu et al., 2022; Kipf, 2016). Several important works consider inverse problems where the quantity inferred lives of a graph description of the geometry. In Garcia Trillos & Sanz-Alonso (2018) the authors prove the convergence of the graph posterior to the continuum posterior subject to some conditions. In Povala et al. (2022) the adjacency structure of a mesh is leveraged to accelerate variational inference inversion. Important works consider solving the PDE forward problem through physics informed graph machine learning(Seo & Liu, 2019; Valencia et al., 2025; Wandel et al., 2020) Used in our comparisons are linear inverse problems with Gaussian process priors, this leads to graph Gaussian process methods (Borovitskiy et al., 2021; Mostowsky et al., 2024). We find it important to mention works in the physics informed literature where (assuming complete or partial knowledge on PDEs, boundary conditions, forcing etc.) one can use physics residuals to learn forward and inverse maps for PDEs (Raissi et al., 2019; Gao et al., 2022; Vadeboncoeur et al., 2023; Xu et al., 2023; Rixner & Koutsourelakis, 2021). However, in this work, we focus solely on methods which do not assume knowledge of the governing PDE, boundary conditions, or forcing and are entirely data-driven in their approach.
@@ 1.2.3 Statistical Autoencoders
Related to inverse problem are latent variable representations of graph data structures. In this category, we focus on the seminal work of the graph VAE (Kipf & Welling, 2016; Nazari et al., 2023). We refer the reader to Appendix C for a more elaborate discussion on the relationship between VAEs and our proposed framework. Related to this are the works (Mylonas et al., 2021; Liu et al., 2019). The proposed work centres around a class of autoencoders dubbed statistical autoencoders, of which there are many variants (Zhao et al., 2019; Kolouri et al., 2019; Turinici, 2021; Tolstikhin et al., 2018). We specifically focus on the MMD (Gretton et al., 2012) variant of the autoencoder. Many other variants could be considered. Another relevant class of autoencoders are conditional autoencoders (Sohn et al., 2015), we will make use of related ideas by conditioning encoder and decoder on geometry.
@@ 1.3 Notation
We denote by $\mathbb{P}_{z}$ a probability measure on the measurable space $(\mathcal{Z},\mathcal{S})$, and by $p_{z}$ the associated density with respect to Lebesgue measure when it exists. Empirical measures are a mixture of Dirac measures. That is, for $D=\{x_{i}\}_{i=1}^{M}$, we have $\widehat{\mathbb{E}}_{x\in D}[\delta_{x}]=\frac{1}{M}\sum_{i=1}^{M}\delta_{x_{i}}$ where $\delta_{x_{i}}(A)=1\text{ if }x_{i}\in A,\text{ and }\delta_{x_{i}}(A)=0,\text{ if }x_{i}\notin A.$ We will sometimes refer to such probability measures as the “densities” $p,q$ to lighten notation. Superscripts in parentheses denote realizations of random variables. For a function $g:\mathcal{Z}\rightarrow\mathcal{U}$, the pushforward $g_{\#}p_{z}$ means $g_{\#}\mathbb{P}_{z}(A)=\mathbb{P}_{z}(g^{-1}(A))$ for all measurable sets $A$. We note, a sample $u^{(n)}\sim g_{\#}p_{z}$ is obtained as $u^{(i)}=g(z^{(i)}),z^{(i)}\sim p_{z}$.
@@ 2.1 Posterior from a Pushforward Prior
Stated simply, in Bayesian inference, when a prior on $\mathcal{U}$ is specified as the pushforward under $g$ (Subsection 1.3) of a latent prior on a space $\mathcal{Z}$, two distributions are equivalent: i) the Bayesian posterior on $\mathcal{U}$, and ii) the pushforward under $g$ of the posterior associated to the latent prior on $\mathcal{Z}$. This result will be explicitly related to the geometric inference setting in Theorem 2.2, showing that with a latent prior constructed via a geometry aware encoder/decoder, we can solve inverse problems on $\mathcal{U}$ with a single common latent representation relevant to any number of in-distribution geometries. All proofs can be found in Appendix A.
@@ 2.2 Geometric Autoencoder Priors for Bayesian Inversion (GABI)
We describe the two-step process of learning a geometry-aware prior and how to use this to solve inverse problems.
@@
Both $E^{\theta}$ (the encoder) and $D^{\psi}$ (the decoder) are graph neural networks with parameter sets $\theta,\psi$ respectively. These two maps imply a geometry conditioned autoencoder. We use $\widehat{\mathbb{E}}_{\mathcal{D}}[\,\cdot\,]$ as a shorthand for $\frac{1}{N}\sum_{(\mathbf{u}_{n},\mathsf{M}_{n})\in\mathcal{D}}[\,\cdot\,]$. Set the parameters of the auto-encoder to
@@
where $p^{\theta}_{\mathbf{z}}=\widehat{\mathbb{E}}_{\mathcal{D}}\,[\delta_{E_{n}^{\theta}(\mathbf{u}_{n})}]$ is the empirical distribution of encoded solutions, $q_{\mathbf{z}}=\mathcal{N}(\bf 0,\mathbf{I})$, $\mathrm{arg\,min}$ is to be approximated with gradient descent, and $\mathsf{d}:\mathcal{P}(\mathsf{Z})\times\mathcal{P}(\mathsf{Z})\rightarrow[0,\infty)$ may be any suitable divergence that is well-posed for comparing empirical measures: maximum mean discrepancy (MMD), energy distance (ED), Wasserstein, or Sliced-Wasserstein distances etc. In this work, we focus on MMD.
@@
$\mathbf{y}_{o}=\mathbf{H}_{o}D^{\psi}_{o}(\mathbf{z})+\bm{\xi}_{o},\quad\bm{\xi}_{o}\sim\mathcal{N}(\mathbf{0},\sigma^{2}{\mathbf{I}}_{o}),$ (4)
where $D^{\psi}_{o}=D^{\psi}(\,\cdot\,;\mathsf{M}_{o})$. We specify our prior distribution over $\mathbf{u}_{o}$, denoted $p_{\mathbf{u}_{o}}$, using the trained decoder; and specify the likelihood, denoted $p^{\psi}_{\mathbf{y}_{o}|\mathbf{z}}$, associated to this decoded observation model (4)
$\centering p_{\mathbf{u}_{o}}^{\psi}:=D^{\psi}_{o\,\#}q_{\mathbf{z}};\quad\quad p^{\psi}_{\mathbf{y}_{o}|\mathbf{z}}=\mathcal{N}(\mathbf{H}_{o}D^{\psi}_{o}(\mathbf{z}),\sigma^{2}\mathbf{I}_{o}).\@add@centering$ (5)
@@
Assuming the relevant measurability of $D^{\psi}_{o}$, the posterior
$p^{\psi}_{\mathbf{u}_{o}|\mathbf{y}_{o}}=D^{\psi}_{o\,\#}p_{\mathbf{z}|\mathbf{y}_{o}}.$ (7)